OpenStudy (anonymous):

You roll a number cube 20 times. The number 4 is rolled 8 times.What is the experimental probability of rolling a 4? how do I figure this out?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

What is the probability of rolling a 4 with one roll of the number cube?

OpenStudy (anonymous):

1 ?

OpenStudy (anonymous):

or 4?

OpenStudy (whpalmer4):

Ah, wait, you don't need that, you're doing empirical probabilities. The empirical probability (or experimental probability) is the ratio of specific events to total number of trials. For example, if you rolled a number cube 100 times, and 80 of those times it came up 6, the empirical probability of rolling a 6 would be 80/100 or 0.8. One might conclude that either more than one face had a 6 on it, or the number cube was rigged.

OpenStudy (anonymous):

ok writing this down just sec.

OpenStudy (anonymous):

trying to understand, sorry

OpenStudy (anonymous):

40% chance?

OpenStudy (whpalmer4):

You could do a trial where you roll your number cube say 100 times, and make a histogram of how many times it comes up with each different number.

OpenStudy (anonymous):

ok historgram I remember those. But is 40% right?

OpenStudy (whpalmer4):

Here's a histogram of a trial I just made up :-)

OpenStudy (anonymous):

thanks I printed it out.

OpenStudy (whpalmer4):

The experimental probability of rolling a 4 would be 10/100, because it came up 10 times in 100 trials. The experimental probability of rolling a 6 would be 5/100, because it came up 5 times in 100 trials.

OpenStudy (whpalmer4):

With a fair number cube, one would expect those bars to be much more equal in height, when enough trials are done.

OpenStudy (anonymous):

Ok so not %40 probability of rolling a four. Writing notes down.

OpenStudy (whpalmer4):

With "small" numbers of trials, that might not be true. For example, the probability of flipping a coin and getting it to land "heads" is 1/2, right? But if you flip the coin only 3 times, you might well manage to flip 3 heads in a row, even though the expected probability of doing so is only 1/2 * 1/2 * 1/2 = 1/8 or 12.5%. Flip the coin 1000 times in a row, and I'm going to be rather unwilling to let you use that coin for making bets with me if it comes up heads 1000 times in a row (the experimental probability of getting a "heads" result there being 1000/1000 = 1, which is much greater than the expected probability of 1/2!)

OpenStudy (whpalmer4):

Well, your experimental probability with your data is 8/20 = 4/10 = 0.4 or 40%. You had a specific event (rolling a 4) which happened 8 times out of 20 trials, so 8/20 = 0.4 or 40%. That's the experimental probability of rolling a 4 with that number cube in that experiment.

OpenStudy (anonymous):

Yeah:)! I got it right. You helped me understand so much. I wrote all the things you said down. Keeping my notes. : ) Thanks you.

OpenStudy (anonymous):

I use to have no problems with probability, but lately I don't what's going on with me. Maybe summer coming.

OpenStudy (whpalmer4):

Start playing dice games with your friends for money — you'll improve your understanding quickly :-)

OpenStudy (anonymous):

fun too! ok thanks!!